2. If rolling two 6-sided dice, find the probability of obtaining a sum of 7.
Answer #1

Total outcomes of rolling two 6-sided dice = 6*6 = 36

total chance to obtaining a sum of 7 : (1, 6), (2, 5), (3, 4), (4, 3), (5, 2), (6, 1) = total 6.

hence probability to obtaining a sum of 7 = 6/36 = 1/6.